# Eliciting and Distinguishing Between Weak and Incomplete Preferences: Theory, Experiment and Computation

### v5

@@ 6         Concluding Remarks
This paper proposes and implements theory-guided methods of data collection and analysis that aim to contribute towards eliciting a decision maker’s possibly weak or incomplete preferences and, where relevant, towards distinguishing between their indifference and indecisiveness parts. On the data-collection side, the paper contributes an incentivized experimental design with multi-valued forced- and free-/non-forced choice treatments. On the data-analytic side, it deploys a model-rich combinatorial-optimization method that allows for recovering an individual’s possibly weak or incomplete preferences from their multi-valued choices in a model-optimal way. This method is a novel extension of the celebrated Houtman and Maks (1985) approach that is fine-tuned to the generally multi-valued nature of the choice data by accounting for the (Jaccard) similarity between the observed and model-optimal choices, and penalizing a subject’s proximity score relative to a model accordingly. We showed how this method is fruitfully applicable on the general model of rational choice with potential indifferences, as well as to two models of incomplete-preference maximization that allow for –and distinguish between– indifference and/or incompleteness. The method is obviously general, however, and in the future can be applied on other models of bounded rationality that predict multi-valued choices.
Despite the relatively large number of decisions, the behaviour of 56% of all subjects in our sample is either perfectly or approximately matched by some simple but richly structured deterministic model of complete or incomplete preference maximization, with uniquely recovered and indifference-exhibiting preferences in the vast majority of cases. Rational choice accounts for the behaviour of over 60% of all subjects in this group (34% of the total), while the two models of incomplete-preference maximization together account for the remaining subjects. Importantly, the optimal preference relation that is recovered conditional on a subject’s best-matching model typically features a non-trivial indifference relation, and therefore highlights the importance of accounting for indifferences in revealed-preference analyses. Finally, exploiting two relevant aspects of the experimental design allows for inferring that a further 10% of all analysed subjects could be thought of as engaging in systematic satisficing behaviour or displaying a robust preference for randomization. This increases to 66% the proportion of individuals whose behaviour can be categorized as following a specific pattern.
Our analysis points to the usefulness of multi-valued and indifference-/incomparability-permitting choice experiments. It also highlights the importance of methodologically pluralistic methods that allow for recovering preferences and/or choice rules by comparing observed behavioural with the predictions made by utility maximization as well as by models of bounded-rational choice or decision heuristics.
Appendices to
“Towards Eliciting Weak or Incomplete Preferences in the Lab: A Model-Rich Approach”
@@ Appendix B Model Goodness-of-Fit Based on the Binary Extension of HM
Undominated Dominant Utility Choice with Choice with Maximization Incomplete Incomplete All Preferences Preferences Forced-Choice treatment ($N=138$) % of subjects with score $=$ 0 3.62% 0% 0% 3.62% % of subjects with score $\leq$ 5 33.33% 2.17% 0% 35.50% % of subjects with score $\leq$ 10 46.38% 8.70% 0% 55.07% Mean/median best score ($\leq 10$) 4.06/4 7.25/8 4.56/4 Minimum score in simulations 25 23 23 Mean/median best-model 1.17/1 1.00/1 1.15/1 preference orderings (score $\leq 10$) Non-Forced-Choice treatment ($N=135$) % of subjects with score $=$ 0 2.22% 0% 5.19% 7.41% % of subjects with score $\leq$ 5 18.51% 2.96% 21.48% 42.96% % of subjects with score $\leq 10$ 24.44% 5.19% 28.15% 57.78% Mean/median best score ($\leq 10$) 4.15/3 5.57/5 3.39/3 3.91/3 Minimum score in simulations 27 26 18 18 Mean/median best-model 1.00/1 1.29/1 1.08/1 1.06/1 preference orderings (score $\leq 10$)
@@
Appendix C Motivation for Strongly Symmetric Menu Collections
@@
Appendix D Choice Probabilities in Simulations
@@
Probability of each Probability of each Probability of non-empty submenu active-choice alternative deferral being chosen being chosen Forced-Choice simulations $\frac{1}{2^{k}-1}$ $\frac{1}{2}\frac{2^{k}}{2^{k}-1}$ NA Non-Forced-Choice simulations $\frac{1}{2^{k}-1}\frac{k}{k+1}$ $\frac{1}{2}\frac{2^{k}}{2^{k}-1}\frac{k}{k+1}$ $\frac{1}{k+1}$

### v6

@@ 6   Concluding Remarks
This paper proposes and implements theory-guided methods of data collection and analysis that aim to contribute towards eliciting a decision maker’s possibly weak or incomplete preferences and, where relevant, towards distinguishing between their indifference and indecisiveness parts. On the data-collection side, the paper contributes an incentivized experimental design with multi-valued forced- and free-choice treatments. On the data-analytic side, it deploys a model-rich combinatorial-optimization method that allows for recovering an individual’s possibly weak or incomplete preferences from their multi-valued choices in a model-optimal way. This method is a novel extension of the celebrated Houtman and Maks (1985) approach that is fine-tuned to the generally multi-valued nature of the choice data by accounting for the (Jaccard) similarity between the observed and model-optimal choices, and penalizing a subject’s proximity score relative to a model accordingly. We showed how this method can be fruitfully applied on the general model of rational choice with potential indifferences, as well as to two models of incomplete-preference maximization that allow for –and distinguish between– indifference and/or incompleteness. The method is obviously general, however, and in the future can be applied on other models of bounded rationality that predict multi-valued choices.
Despite the relatively large number of decisions, the behaviour of 56% of all subjects in our sample is either perfectly or approximately matched by some simple but richly structured deterministic model of complete or incomplete preference maximization, with uniquely recovered and indifference-exhibiting preferences in the vast majority of cases. Rational choice accounts for the behaviour of over 60% of all subjects in this group (34% of the total), while the two models of incomplete-preference maximization together account for the remaining subjects. Importantly, the optimal preference relation that is recovered conditional on a subject’s best-matching model typically features a non-trivial indifference relation that may encompass up to 19% of all possible comparisons, and therefore highlights the importance of accounting for indifferences in revealed-preference analyses.
Our analysis points to the usefulness of multi-valued and indifference-/incomparability-permitting choice experiments. It also highlights the importance of methodologically pluralistic methods that allow for recovering preferences and/or choice rules by comparing observed behavioural with the predictions made by utility maximization as well as by models of bounded-rational choice.
@@ Appendix B Model Goodness-of-Fit Based on the Binary Extension of HM
Undominated Dominant Utility Choice with Choice with Maximization Incomplete Incomplete All Preferences Preferences Forced-Choice treatment ($N=138$) % of subjects with score == 0 3.62% 0% 0% 3.62% % of subjects with score $\leq$ 5 33.33% 2.17% 0% 35.50% % of subjects with score $\leq$ 10 46.38% 8.70% 0% 55.07% Mean/median best score ($\leq 10$) 4.06/4 7.25/8 4.56/4 Minimum score in simulations 25 23 23 Mean/median best-model 1.17/1 1.00/1 1.15/1 preference orderings (score $\leq 10$) Non-Forced-Choice treatment ($N=135$) % of subjects with score == 0 2.22% 0% 5.19% 7.41% % of subjects with score $\leq$ 5 18.51% 2.96% 21.48% 42.96% % of subjects with score $\leq 10$ 24.44% 5.19% 28.15% 57.78% Mean/median best score ($\leq 10$) 4.15/3 5.57/5 3.39/3 3.91/3 Minimum score in simulations 27 26 18 18 Mean/median best-model 1.00/1 1.29/1 1.08/1 1.06/1 preference orderings (score $\leq 10$)
@@
Appendix C Analysis of Other Decision Models & Heuristics
Although not constructed so as to enable a thorough investigation of such behaviours, utilizing two important aspects of the experimental design and its implementation allows for inferring that an additional 10% of the analysed subjects could be thought of as exhibiting a systematic satisficing behaviour (Simon, 1956; Caplin, Dean, and Martin, 2011; Reutskaja, Nagel, Camerer, and Rangel, 2011) or a preference for randomization (Agranov and Ortoleva, 2017, 2023; Dwenger, Kübler, and Weizsäcker, 2018; Cettolin and Riedl, 2019; Agranov, Healy, and Nielsen, 2022). Additionally, while the experiment was not designed to test for this either, we also investigated the potential presence of a “preference for flexibility” over menus (Kreps, 1979; Dekel, Lipman, and Rustichini, 2001). We did so by entertaining the possibility that some subjects might have “metavisualised” the actual experimental task of choosing potentially many alternatives from each menu and responded to it by choosing from the collection of all menus that are derivable from that menu. Reassuringly, this analysis does not detect any notable patterns pointing towards a systematic such “metavisualised” preference for flexibility. Details of these additional analyses are presented in this appendix.
C.1   Satisficing
In addition, we compute and study the average position of each subject’s chosen item(s) in the 50 menus’ list orderings because, intuitively, an unusually low average value of this metric could also be indicative of satisficing behaviour for the subject in question. We find that for an additional 7 and 5 (4.4%) subjects in the Forced- and Free-Choice treatments, respectively, who were not classified as approximately explainable by one of the three deterministic models that were discussed above, the average positions of their chosen item(s) in the 50 menus’ list orderings were strictly below the 2.5% percentile cut-off value of 1.84 that is derived from simulated uniform-random forced choices. These two distinct analyses therefore suggest that a total 21 subjects (7.7%) who were not model-classified in the main part of the paper might be thought of as exhibiting a systematic satisficing behaviour.
(a) Forced-Choice treatment
Distribution range: $[1.15,2.25]$
(b) Free-Choice treatment
Distribution range: $[1.74,2.32]$
Note: $p=0.009$; two-sided Mann-Whitney $U$ test.
C.2   Preference for Randomization
C.3   Menu Meta-Visualization and Menu Preferences for Flexibility
The leading theories of menu preferences in the literature portray decision makers as exhibiting either a preference for flexibility (Kreps, 1979; Dekel, Lipman, and Rustichini, 2001) or a preference for commitment (Gul and Pesendorfer, 2001, 2005). [^13] Toussaert (2018) reports on an experimental design that was explicitly constructed to test for the presence of such preferences. The former preference is manifested at a menu $A$ if $B$ is preferred to $A$ whenever $B\supset A$. By contrast, a preference for commitment in the Gul-Pesendorfer theory of temptation and self-control is formalized with the Set Betweeness axiom whereby “$A$ is preferred to BB” implies “$A$ is preferred to $A\cup B$ is preferred to BB”.
However, the basic aggregate-level analysis in the main part of the paper readily suggests that few –if any– subjects in our sample of eligible subjects are likely to exhibit a systematic preference for flexibility. [^14] Four subjects in the original sample could indeed be thought of as showing –trivially– such a consistent preference for flexibility throughout the experiment by always choosing all alternatives at every menu. As noted in the main part of the paper, those subjects were excluded from the subsequent analysis because their behaviour was completely uninformative about their preferences over gift-card pairs. This point is validated in Figure 14, which depicts the subjects’ cumulative densities of flexibility violations. More specifically, 5 and 12 are the smallest numbers of menus where this principle was violated, in each case by a single subject, while the means/medians in the two treatments stand at 45.8/48 (Forced Choice) and 46.6/49 (Free Choice). Therefore, in line with the caveats expressed at the beginning of this subsection, looking at our experimental data through this lens does not seem to add any notable new insights into subjects’ behaviour. This provides some further reassurance toward the validity of the results presented earlier in this section.
Appendix D Motivation for Strongly Symmetric Menu Collections
@@
Appendix E Choice Probabilities in Simulations
@@
Probability of each Probability of each Probability of non-empty submenu active-choice alternative deferral being chosen being chosen Forced-Choice simulations $\frac{1}{2^{k}-1}$ $\frac{1}{2}\frac{2^{k}}{2^{k}-1}$ NA Free-Choice simulations $\frac{1}{2^{k}-1}\frac{k}{k+1}$ $\frac{1}{2}\frac{2^{k}}{2^{k}-1}\frac{k}{k+1}$ $\frac{1}{k+1}$
@@
Appendix F Graphs of Model-Optimal Revealed Preference Orderings
Note: arrows point to dominated options/indifference classes
F.1   Rational Choice/Utility Maximization
(a) ID=1140, JHM=0,
1 optimal relation
(b) ID=2582, JHM=0,
1 optimal relation
(c) ID=3843, JHM=0,
1 optimal relation
(d) ID=7228, JHM=0,
1 optimal relation
(a) ID=7538, JHM=0,
1 optimal relation
(b) ID=9139, JHM=0,
1 optimal relation
(c) ID=9425, JHM=0,
1 optimal relation
(d) ID=9507, JHM=0,
1 optimal relation
(a) ID=9946, JHM=0.333,
1 optimal relation
(b) ID=1411, JHM=0.5,
1 optimal relation
(c) ID=1967, JHM=0.5,
1 optimal relation
(d) ID=2895, JHM=0.5,
1 optimal relation
(a) ID=6861, JHM=0.5,
1 optimal relation
(b) ID=8360, JHM=0.5,
1 optimal relation
(c) ID=6969, JHM=0.5,
1 optimal relation
(d) ID=8712, JHM=0.5,
1 optimal relation
(a) ID=4430, JHM=0.5,
1 optimal relation
(b) ID=4540, JHM=0.5,
1 optimal relation
(c) ID=3797, JHM=0.75,
1 optimal relation
(d) ID=6679, JHM=0.833,
1 optimal relation
(a) ID=8722, JHM=0.833,
1 optimal relation
(b) ID=4542, JHM=1,
1 optimal relation
Refer to captionRefer to caption
(c) ID=6890, JHM=1,
3 optimal relations
(d) ID=7081, JHM=1,
1 optimal relation
(a) ID=3622, JHM=1,
1 optimal relation
(b) ID=3826, JHM=1,
1 optimal relation
(c) ID=6778, JHM=1,
1 optimal relation
(d) ID=7464, JHM=1,
1 optimal relation
(a) ID=7752, JHM=1,
1 optimal relation
(b) ID=3986, JHM=1.167,
1 optimal relation
(c) ID=6112, JHM=1.167,
1 optimal relation
(d) ID=3452, Score=1.333,
1 optimal relation
(a) ID=8855, JHM=1.333,
1 optimal relation
(b) ID=1230, JHM=1.333,
1 optimal relation
Refer to captionRefer to caption
(c) ID=5134, JHM=1.5,
3 optimal relations
(d) ID=8553, JHM=1.5,
1 optimal relation
(a) ID=1827, JHM=1.5,
1 optimal relation
(b) ID=3377, JHM=1.5,
1 optimal relation
(c) ID=3473, JHM=1.5,
1 optimal relation
(d) ID=5337, JHM=1.5,
1 optimal relation
(a) ID=8773, JHM=1.5,
1 optimal relation
(b) ID=7677, JHM=1.5,
1 optimal relation
(c) ID=1557, Score=1.667, 1 optimal relation
(d) ID=9039, JHM=1.667,
1 optimal relation
(a) ID=4753, JHM=1.667,
1 optimal relation
Refer to captionRefer to caption
(b) ID=5061, JHM=1.833,
3 optimal relations
(c) ID=6923, JHM=1.833,
1 optimal relation
(d) ID=2952, JHM=1.833,
1 optimal relation
(a) ID=9766, JHM=2,
1 optimal relation
(b) ID=9934, JHM=2,
1 optimal relation
Refer to captionRefer to caption
(c) ID=1068, JHM=2,
2 optimal relations
(d) ID=1609, JHM=2,
1 optimal relation
(a) ID=6222, JHM=2,
1 optimal relation
Refer to captionRefer to caption
(b) ID=8468, JHM=2,
2 optimal relations
(c) ID=8989, JHM=2,
1 optimal relation
(d) ID=5820, JHM=2.167,
1 optimal relation
(a) ID=4597, JHM=2.25,
1 optimal relation
(b) ID=4261, JHM=2.25,
1 optimal relation
(c) ID=9460, JHM=2.417,
1 optimal relation
(d) ID=5878, JHM=2.5,
1 optimal relation
Refer to captionRefer to caption
(a) ID=2437, JHM=2.5,
3 optimal relations
Refer to captionRefer to caption
(b) ID=4152, JHM=2.5,
2 optimal relations
(c) ID=8892, JHM=2.667,
1 optimal relation
(d) ID=5323, JHM=3,
1 optimal relation
(a) ID=7996, JHM=3,
1 optimal relations
(b) ID=2155, JHM=3,
1 optimal relation
(c) ID=3788, JHM=3,
1 optimal relation
(d) ID=4924, JHM=3,
1 optimal relation
(a) ID=8732, JHM=3,
1 optimal relation
(b) ID=9899, JHM=3,
1 optimal relation
Refer to captionRefer to caption
(c) ID=9292, JHM=3.167,
2 optimal relations
(d) ID=4652, JHM=3.167,
1 optimal relation
(a) ID=6384, JHM=3.167,
1 optimal relation
(b) ID=7215, JHM=3.5,
1 optimal relation
(c) ID=9134, JHM=3.583,
1 optimal relation
(d) ID=2864, JHM=3.667,
1 optimal relation
(a) ID=9985, JHM=3.667,
1 optimal relation
Refer to captionRefer to caption
(b) ID=1873, JHM=3.667,
3 optimal relations
(c) ID=4587, JHM=3.75,
1 optimal relation
(d) ID=5646, JHM=4,
1 optimal relation
Refer to captionRefer to caption
(a) ID=2753, JHM=4,
2 optimal relations
(b) ID=1727, JHM=4,
1 optimal relation
(c) ID=6715, JHM=4.5,
1 optimal relation
(d) ID=6657, JHM=4.833,
1 optimal relation
Refer to captionRefer to caption
(a) ID=4285, JHM=4.833,
2 optimal relations
(b) ID=6535, JHM=4.833,
1 optimal relation
(c) ID=7735, JHM=4.833,
1 optimal relation
(a) ID=8542, JHM=4.833,
1 optimal relation
(b) ID=9294, JHM=4.917,
1 optimal relation
(c) ID=7763, JHM=5,
1 optimal relation
(d) ID=7332, JHM=5,
1 optimal relation
F.2   Undominated Choice with Incomplete Preferences
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 39: ID=1889, JHM=0.667, 15 optimal relations
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 40: ID=9001, JHM=1, 15 optimal relations
(a) ID=8945, JHM=1.333,
1 optimal relation
Refer to captionRefer to caption
Refer to captionRefer to caption
(b) ID=6948, JHM=1.5, 4 optimal relations
(a) ID=8721, JHM=1.5,
1 optimal relation
Refer to captionRefer to caption
Refer to captionRefer to caption
(b) ID=6521, JHM=2.083, 4 optimal relations
Refer to captionRefer to caption
(a) ID=6140, JHM=2.167, 2 optimal relations
Refer to captionRefer to caption
(b) ID=2432, JHM=3, 2 optimal relations
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 44: ID=7017, JHM=3.083, 10 optimal relations
Refer to captionRefer to caption
(a) ID=9292, JHM=3.167,
2 optimal relations
(b) ID=9222, JHM=3.333,
1 optimal relation
Refer to captionRefer to caption
(c) ID=8075, JHM=3.5,
2 optimal relations
Refer to captionRefer to caption
(a) ID=8719, JHM=3.667,
2 optimal relations
(b) ID=3574, JHM=4,
1 optimal relation
(c) ID=7423, JHM=4,
1 optimal relation
(a) ID=9819, JHM=4,
1 optimal relation
Refer to captionRefer to caption
(b) ID=5049, JHM=4.167,
2 optimal relations
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 48: ID=5283, JHM=4.167, 5 optimal relations
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 49: ID=8049, JHM=4.25, 12 optimal relations
(a) ID=2038, JHM=4.667,
1 optimal relation
Refer to captionRefer to caption
(b) ID=2375, JHM=4.667,
2 optimal relations
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 51: ID=1308, JHM=4.833, 15 optimal relations
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 52: ID=3645, JHM=4.833, 4 optimal relations
F.3   Dominant Choice with Incomplete Preferences
(a) ID=1716, JHM=0,
1 optimal relation
(b) ID=3487, JHM=0,
1 optimal relation
(c) ID=3746, JHM=0,
1 optimal relation
(a) ID=4486, JHM=0,
1 optimal relation
(b) ID=5649, JHM=0,
1 optimal relation
(c) ID=5813, JHM=0,
1 optimal relation
(a) ID=9142, JHM=0,
1 optimal relation
(b) ID=2558, JHM=0.333,
1 optimal relation
(c) ID=7128, JHM=0.5,
1 optimal relation
(a) ID=8407, JHM=0.5,
1 optimal relation
(b) ID=4112, JHM=1,
1 optimal relation
(c) ID=5317, JHM=1,
1 optimal relation
(a) ID=8555, JHM=1,
1 optimal relation
(b) ID=2882, JHM=1.5,
1 optimal relation
(c) ID=5591, JHM=1.5,
1 optimal relation
(a) ID=2855, JHM=2,
1 optimal relation
(b) ID=3116, JHM=2,
1 optimal relation
(c) ID=7286, JHM=2,
1 optimal relation
(a) ID=7621, JHM=2,
1 optimal relation
(b) ID=9072, JHM=2,
1 optimal relation
(c) ID=9650, JHM=2.083,
1 optimal relation
(a) ID=1124, JHM=2.167,
1 optimal relation
(b) ID=5064, JHM=2.333,
1 optimal relation
(c) ID=5860, JHM=2.333,
1 optimal relation
(a) ID=7481, JHM=2.333,
1 optimal relation
(b) ID=4155, JHM=2.5,
1 optimal relation
(c) ID=8801, JHM=2.5,
1 optimal relation
(d) ID=5131, JHM=3,
1 optimal relation
(e) ID=5440, JHM=3,
1 optimal relation
(a) ID=6055, JHM=3,
1 optimal relation
(b) ID=7869, JHM=3.5,
1 optimal relation
(c) ID=8133, JHM=3.5,
1 optimal relation
Refer to captionRefer to caption
(a) ID=8781, JHM=3.667, 3 optimal relations
(b) ID=9305, JHM=4,
1 optimal relation
(a) ID=5580, JHM=4.5,
1 optimal relation
(b) ID=7071, JHM=4.833,
1 optimal relation
(c) ID=7577, JHM=5,
1 optimal relation